OpenStudy (anonymous):

Square root: 2x-6=4 solve the radical equation for x,

OpenStudy (anonymous):

is it \[\sqrt{2x - 6} = 4\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

square both sides and tell me what u get..... then i will guide u further...

OpenStudy (anonymous):

heydave??????

OpenStudy (anonymous):

if u want to progress, then try to work out the problem with our help... so square both sides and tell me what you get...

OpenStudy (anonymous):

squaring both sides we get 2x - 6 = 16 2x = 16 + 6 x = 22/2 x = 11